# A Neurosymbolic Fast and Slow Architecture for Graph Coloring

### v1

@@ Discussion
This study broadens the scope of neurosymbolic systems by applying them to graph coloring problems. It demonstrates that integrating LLMs with traditional symbolic solvers, guided by metacognitive governance, significantly enhances solutions’ accuracy and efficiency. Unlike previous applications that primarily leveraged LLMs for sequential decision-making tasks (Valmeekam et al. 2022; Stechly, Valmeekam, and Kambhampati 2024), SOFAI-v2 directly addresses their constraint adherence limitations through iterative feedback and strategic use of symbolic reasoning. This approach aligns with cognitive decision-making theories and shows a better success rate and time efficiency, illustrating the practical benefits of hybrid architectures in complex constraint satisfaction problems.
Conclusion
@@
We introduced SOFAI-v2, a neurosymbolic architecture that integrates a fast, episodic memory-based LLM (S1) with a slow, deliberative symbolic solver (S2), enhanced by real-time metacognitive governance. Our empirical evaluations across various graph coloring problems confirm that SOFAI-v2 outperforms traditional symbolic solvers and its predecessor, SOFAI-v1, in both success rates and time efficiency, especially within complex problem configurations.
SOFAI-v2 employs iterative metacognitive feedback and episodic memory-based iterations to enable S1 to refine its response. This effectively overcomes the limitations of LLMs in constraint adherance of CSPs. By leveraging episodic memory and adaptive feedback, SOFAI-v2 significantly enhances the accuracy and efficiency of solving CSPs, demonstrating the potential of cognitive-inspired AI architectures in managing complex challenges. Future work will scale SOFAI-v2 to tackle larger and more varied problem sets, explore different graph models, and apply the architecture to other types of CSPs and sequential decision-making tasks. Enhancing the metacognitive components with advanced learning algorithms and incorporating alternative neural models are anticipated to boost its adaptability and efficiency further, contributing to developing robust, scalable solutions for intricate AI applications.
Appendix A Prompt, Example, and Feedback Examples
@@
Please provide the color assignment for the new problem to solve, or respond with ”NOT SOLVABLE” if it cannot be solved.
Figure 8: LLM prompt template for $m(0,100)$ and $m(50,50)$ for the graph coloring problem without using episodic memory.
@@
This example is provided to demonstrate a previously successful coloring strategy for a similar subgraph, which may be helpful in solving the current problem. Please provide the color assignment for the new problem to solve directly below, or respond with ”NOT SOLVABLE” if it cannot be solved.
@@
Appendix B Results - Success Rate
@@
Appendix C Results - Average Time Take
This section explores the time taken by different solvers to solve graph coloring problems across different problem configurations ($m$). Figure 16, Figure 17, and Figure 18 demonstrate how each solver—System 1, System 2, SOFAI-v1, and SOFAI-v2—performs across varying edge probabilities ($p$). These graphs show the time efficiency of SOFAI-v2 compared to System 2 and SOFAI-v1.

### v2

@@ 5.6 Implications and Comparisons with Prior Work
This study broadens the scope of neurosymbolic systems by applying them to graph coloring problems. It demonstrates that integrating LLMs with traditional symbolic solvers, guided by metacognitive governance, enhances accuracy and efficiency. Unlike previous applications that primarily leveraged LLMs for sequential decision-making tasks (Valmeekam et al., 2022; Stechly et al., 2024), SOFAI-v2 directly addresses constraint-adherence limitations through iterative feedback and strategic use of symbolic reasoning. Viewed through cognitive control, S1 provides routine proposals, while MC and S2 supply supervisory/controlled processing when conflicts persist, consistent with contention scheduling and supervisory attention (Norman & Shallice, 1986; Cooper & Shallice, 2000). This neurosymbolic framing clarifies why SOFAI-v2 improves success rate and time efficiency on constraint satisfaction problems.
6 Conclusion and Future Work
@@
We introduced SOFAI-v2, a neurosymbolic architecture that integrates a fast, episodic memory-based LLM (S1) with a slow, deliberative symbolic solver (S2), enhanced by real-time metacognitive governance. Our empirical evaluations across various graph coloring problems confirm that SOFAI-v2 outperforms traditional symbolic solvers and its predecessor, SOFAI-v1, in both success rates and time efficiency, especially within complex problem configurations. SOFAI-v2 employs iterative metacognitive feedback and episodic memory-based iterations to enable S1 to refine its response. This effectively overcomes the limitations of LLMs in constraint adherence for CSPs. By leveraging episodic memory and adaptive feedback, SOFAI-v2 improves the accuracy and efficiency of solving CSPs, indicating the value of cognitively informed, neurosymbolic architectures.
The observed improvements are consistent with dual-process theories: S1 generates a fast, potential solution; the metacognitive controller checks it for constraint adherence and uncertainty, requests targeted revisions when needed, and escalates to S2 only when necessary; S2 then performs slow, symbolic search as invoked. Episodic retrieval provides additional context from prior instances to support S1’s proposal and the controller’s checks. In this arrangement, SOFAI-v2 realizes a control loop that finds a trade-off between speed and reliability in a resource-constrained setting.
Future work will scale SOFAI-v2 to larger and more varied problem sets, explore different graph models, and apply the architecture to other CSPs and sequential decision-making tasks. We will also investigate learning-based MC policies and richer memory retrieval, and examine correspondences to cognitive-control signals (e.g., conflict costs) that could enable tighter links to computational cognitive neuroscience. Concretely, we will design human-aligned probes (e.g., speed–accuracy manipulations, conflict-cost perturbations, and memory cueing/interference) to test whether SOFAI-v2’s control dynamics mirror classic findings on supervisory attention and conflict-driven effort allocation.
Appendix
7 Modified DSATUR Algorithm with Backtracking
Given the graph coloring problem $(G,C,k)$, with an undirected graph $G=(V,E)$, a set of available colors $C={c_{1},c_{2},\dots,c_{k}}$, and the maximum allowed number of colors $k$, the modified DSATUR algorithm systematically explores assignments using backtracking. Below is a detailed, mathematically rigorous description:
Definitions:
  • Saturation degree $d_{sat}(v)$ for a vertex $v\in V$ is the cardinality of the set of distinct colors assigned to its adjacent vertices:
  • Vertex degree $d(v)$ is the number of edges incident to $v$.
Algorithm Steps:
  1. 1.
    Initialization: Assign an initial coloring $f(v)=\emptyset$ for all $v\in V$ (all vertices uncolored).
  2. 2.
    Compute Saturation Degree: For each uncolored vertex $v$, compute its saturation degree $d_{sat}(v)$.
  3. 3.
    Select Vertex: Identify vertex $v^{*}$ that maximizes $d_{sat}(v)$. If multiple vertices have equal highest saturation degree, select the one with the highest vertex degree $d(v)$.
  4. 4.
    Color Assignment: Assign to $v^{*}$ the smallest color $c\in C$ that satisfies the condition:
    $f(v^{*})\neq f(u),\quad\forall u\text{ adjacent to }v^{*}$
  5. 5.
    Backtracking: If no valid color can be assigned, recursively backtrack:
    • Undo the previous assignment and select the next smallest available color.
    • Continue this recursive search until a valid color assignment is found or all possibilities are exhausted.
  6. 6.
    Termination: Repeat steps 2-5 until all vertices are successfully colored, or conclusively determine no valid coloring exists under the provided constraint of $k$ colors.
This rigorous exploration via backtracking guarantees the determination of a feasible coloring if one exists within the constraint of using no more than $k$ colors.
8 Algorithm SOFAI-v2
Algorithm Description.
Given a graph $G=(V,E)$ and a color bound $k$, SOFAI-v2 runs an S1–MC loop for at most $T$ iterations. In each iteration, S1 is prompted with the graph, $k$, and the accumulated history. S1 proposes a coloring assignment, which is checked for feasibility. If valid, the algorithm returns the assignment. If invalid, MC extracts the conflicts, generates minimal checkable feedback, and appends this to the history and trend signals. If repeated patterns or the iteration limit is reached, MC escalates to S2. S2 runs a DSATUR backtracking solver for the fixed-$k$ decision problem. If a valid coloring is found, it is returned and stored in memory; otherwise, the algorithm reports “NOT SOLVABLE”.
Algorithm 1 SOFAI-v2 for Graph Coloring (Decision, fixed $k$)
1:Undirected graph $G=(V,E)$, color limit $k$, max S1 iterations $T$, memory $\mathcal{M}$
2:Coloring assignment or “NOT SOLVABLE”
[…]
6:for $t=1$ to $T$ do
7:  $prompt\leftarrow\text{GenerateGCPrompt}(G,k,seed,history)$
8:  $resp\leftarrow\text{CallS1}(prompt)$
9:  $assign\leftarrow\text{ParseColoring}(resp)$
10:  $(ok,viol)\leftarrow\text{CheckFeasibility}(G,k,assign)$
11:  if $ok=\text{true}$ then
12:   return assignassign
13:  end if
14:  $C\leftarrow\text{ExtractConflicts}(viol)$
15:  $H\leftarrow\text{InducedSubgraph}(G,C)$
16:  $f\leftarrow\text{FormatFeedback}(H,C,assign)$
17:  $history\leftarrow history\cup\{(assign,f)\}$
18:  $trends\leftarrow trends\cup\{\text{FeedbackSignature}(f)\}$
19:  if $\text{EscalationRule}(trends,t)$ then
20:   break $\triangleright$ escalate to S2
21:  end if
22:  $seed\leftarrow\text{UpdateSeed}(assign,f)$
23:end for
24:
25:// S2 fallback (decision, fixed $k$)
[…]
27:if $status=\text{SAT}$ then
28:  $\text{UpdateMemory}(\mathcal{M},G,k,assign)$
29:  return assignassign
30:else
31:  return “NOT SOLVABLE”
32:end if
9 Prompt, Example, and Feedback Examples
@@
Please provide the color assignment for the new problem to solve, or respond with "NOT SOLVABLE" if it cannot be solved.
Figure 7: LLM prompt template for the graph coloring decision problem without using episodic memory.
@@
This example is provided to demonstrate a previously successful coloring strategy for a similar subgraph, which may be helpful in solving the current problem. Please provide the color assignment for the new problem to solve directly below, or respond with "NOT SOLVABLE" if it cannot be solved.
@@
10 Results - Average time Solver Comparison
Table 4 reports the average time taken by different solver configurations (S1, S2, SOFAI_v1, and SOFAI-v2) under a fixed time limit of 200 seconds. These results are obtained from the same experiment used to evaluate success rates in Section 5.4. Instances not solved within the allotted 200 seconds are excluded from the average time calculation. Cells with a dash (--) indicate that the solver failed to solve any instance in that configuration within the time limit, resulting in no valid average time. These values help assess not only whether solvers succeed but also how quickly they do so under realistic constraints.
Table 4: Average time taken (in seconds) by solvers under a 200-second time limit. Results are reported for each graph size and solvability configuration. Missing values (dashes) indicate that no problem was solved within the time limit for that solver and configuration. These results correspond to the same experiments used in the success rate analysis in RQ1-Section 5.4.
Graph Size $m(100,0)$ (s) $m(0,100)$ (s) $m(50,50)$ (s)
5 3.23 0.001 3.23 3.92 3.13 0.001 3.13 3.77 3.25 0.001 3.23 3.92
10 3.60 0.19 3.75 3.21 3.80 0.17 3.90 3.07 3.54 0.22 3.85 3.74
15 3.80 54.25 32.90 31.21 3.85 60.36 34.95 28.21 3.21 56.32 53.92 33.92
20 3.85 83.22 54.95 48.72 3.90 99.06 55.28 42.31 3.87 108.9 71.98 55.63
30 - - - - 3.92 - 3.94 19.92 3.92 - 3.97 24.23
40 - - - - 3.95 - 3.96 22.94 3.95 - 3.99 23.25
50 - - - - 3.96 - 3.97 21.76 3.97 - 3.97 24.45
11 Results - Success Rate
@@
For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.
@@
12 Results - Average Time Taken
This section explores the time different solvers take to solve graph coloring problems across different problem configurations ($m$). Figure 15, Figure 16, and Figure 17 demonstrate how each solver—System 1, System 2, SOFAI-v1, and SOFAI-v2—performs across varying edge probabilities ($p$). These graphs show the superiority in time efficiency of SOFAI-v2 compared to System 2 and SOFAI-v1.
For the above results we set a fixed time limit of 200 seconds per problem. If a solver does not finish within this time, it is counted as a failure.